### v1

# Logical-CommonsenseQA: A Benchmark for Logical Commonsense Reasoning

Abstract
Commonsense reasoning often involves evaluating multiple plausible interpretations rather than selecting a single atomic answer, yet most benchmarks rely on single-label evaluation, obscuring whether statements are jointly plausible, mutually exclusive, or jointly implausible. We introduce Logical-CommonsenseQA, a benchmark that reframes commonsense reasoning as logical composition over pairs of atomic statements using plausibility-level operators (AND, OR, NEITHER/NOR). Evaluating instruction-tuned, reasoning-specialized, and fine-tuned models under zero-shot, few-shot, and chain-of-thought prompting, we find that while models perform reasonably on conjunctive and moderately on disjunctive reasoning, performance degrades sharply on negation-based questions. Logical-CommonsenseQA exposes fundamental reasoning limitations and provides a controlled framework for advancing compositional commonsense reasoning.
Logical-CommonsenseQA:
A Benchmark for Logical Commonsense Reasoning
Obed Junias University of Colorado Boulder obed.junias@colorado.edu          Maria Leonor Pacheco University of Colorado Boulder maria.pacheco@colorado.edu
1 Introduction
Commonsense reasoning is central to human cognition and a long-standing challenge in artificial intelligence and natural language understanding. Developmental studies suggest it emerges from intuitive, experience-based frameworks in early childhood and becomes more abstract and socially grounded through embodied experience, social interaction, cultural transmission, and language Spelke et al. (1992); Spelke and Kinzler (2007); Gopnik (2004); Meltzoff et al. (2009). In contrast, large language models (LLMs) acquire “commonsense-like” knowledge indirectly from large-scale text corpora via objectives such as next-token prediction, rather than through grounded interaction or social learning Bender and Koller (2020). Despite this, LLMs exhibit emergent behaviors resembling human intuitive reasoning, including inferring intentions and anticipating socially plausible actions Sap et al. (2019a); Bisk et al. (2020). However, this competence is brittle, as models fail under minimal semantic perturbations, struggle with logical consistency, and rely on statistical shortcuts rather than genuine reasoning Ettinger (2020); Geiger et al. (2020); McCoy et al. (2019); Niven and Kao (2019). To improve the evaluation of commonsense reasoning, we introduce Logical-CommonsenseQA, a benchmark that reframes the task as a logical composition over multiple plausible answers, requiring models to reason about relationships between plausible statements rather than select a single response.
Early approaches to commonsense reasoning relied on symbolic knowledge bases Lenat et al. (1990); Liu and Singh (2004), which were expressive but limited in scalability and coverage. More recent benchmarks shift the evaluation towards the ability of neural models to infer plausible answers from text Talmor et al. (2019); Sap et al. (2019b, a); Zellers et al. (2019); Bisk et al. (2020); Sakaguchi et al. (2021). However, these benchmarks reduce commonsense reasoning to single-answer prediction, despite the fact that many real-world situations admit multiple plausible interpretations Min et al. (2020). For example, in the question “The fox walked from the city into the forest—what was it looking for?” both “natural habitat” and “shelter from disturbances” are plausible answers. Restricting the evaluation to a single atomic choice removes this natural ambiguity and obscures whether models recognize jointly plausible, mutually exclusive, or jointly implausible interpretations.
Beyond linguistic ambiguity, commonsense is both socially grounded and compositional. People judge commonsense statements not only by personal belief but also by estimating whether others would share that belief Whiting and Watts (2024), implying that commonsense reflects a distribution over plausible interpretations rather than a single canonical truth. At the same time, human reasoning integrates multiple alternatives through compositional inference, recognizing when statements are jointly plausible (AND), partially plausible (OR), or jointly implausible (NEITHER/NOR). These relational judgments are central to everyday reasoning about intentions, plans, and social norms, yet current benchmarks fail to evaluate whether models can reason over such socially grounded and compositional plausibility relationships.
In Logical-CommonsenseQA, each instance pairs two atomic answers under one of three symbolic relations—AND, OR, or NEITHER/NOR. These relations are not strict propositional operators, but plausibility-level composition constructs indicating joint, partial, or joint implausibility. This formulation preserves the multiple-choice question (MCQ) format of modern benchmarks while explicitly modeling ambiguity and compositional reasoning. We evaluate instruction-tuned, reasoning-specialized, and fine-tuned LLMs under zero-shot, few-shot, and chain-of-thought prompting. Results show that while models perform reasonably well on conjunctive and moderately on disjunctive reasoning, performance degrades sharply on negation-based compositions, revealing persistent gaps in the commonsense reasoning capabilities of state-of-the-art systems.
2 Related Work
Commonsense reasoning has been widely studied through symbolic knowledge bases, neural models, and benchmark-driven evaluation. Early symbolic approaches such as Cyc Lenat et al. (1990) and ConceptNet Liu and Singh (2004) enabled structured inference but suffered from limited coverage and scalability. More recent benchmarks, including CommonsenseQA Talmor et al. (2019), SocialIQA Sap et al. (2019b), ATOMIC Sap et al. (2019a), PIQA Bisk et al. (2020), HellaSWAG Zellers et al. (2019), and WinoGrande Sakaguchi et al. (2021), shift evaluation towards the ability of neural models to infer plausible responses from text. However, these benchmarks adopt a single atomic-answer format that fails to capture the inherent ambiguity of many real-world commonsense questions.
Prior work has emphasized that ambiguity and human disagreement are intrinsic to commonsense reasoning. AmbigQA Min et al. (2020) and ProtoQA Boratko et al. (2020) show that many questions naturally admit multiple valid answers, motivating the need to evaluate systems beyond single gold answers. Separately, logical reasoning benchmarks such as LogiQA Liu et al. (2020), ReClor Yu et al. (2020), and COM2 Fang et al. (2024) evaluate deductive inference using formal logical operators, but target logical validity rather than commonsense plausibility. Complementary work in the social sciences argues that commonsense is socially grounded, reflecting shared expectations rather than objective truth Whiting and Watts (2024).
Our work differs from these lines of research by introducing Logical-CommonsenseQA, a benchmark that evaluates commonsense reasoning as compositional plausibility. Instead of open-ended multi-answer annotation or formal deductive logic, Logical-CommonsenseQA preserves the MCQ format of modern commonsense benchmarks while explicitly encoding joint plausibility, partial plausibility, and joint implausibility through symbolic composition over human-validated atomic answers.
Operator Interpretation Example
$\boldsymbol{a\;\mathtt{AND}\;b}$ Both statements independently plausible Q: Sammy wanted to go to where the people were. Where might he go?
A: local events AND social venues
$\boldsymbol{a\;\mathtt{OR}\;b}$ At least one statement plausible Q: Sammy wanted to go to where the people were. Where might he go?
A: local events OR empty parks
$\boldsymbol{\mathtt{NEITHER}\;a\;\mathtt{NOR}\;b}$ Neither statement plausible Q: Sammy wanted to go to where the people were. Where might he go?
A: NEITHER quiet retreats NOR empty parks
Table 1: Composition operators used in Logical-CommonsenseQA to express plausibility relations.
3 The Logical Commonsense Dataset
We construct Logical-CommonsenseQA to evaluate compositional, multi-answer commonsense reasoning by extending CommonsenseQA Talmor et al. (2019), a benchmark of 12,247 MCQs with one correct answer and four distractors that probes commonsense relations such as cause–effect, purpose, and spatial association. Logical-CommonsenseQA reformulates these single-answer questions into logically composed items that explicitly model relationships among plausible alternatives while preserving full compatibility with the MCQ format.
Construction Pipeline.
The dataset is built using a three-stage pipeline that integrates neural generation with deterministic symbolic composition, converting atomic answers into logically structured multi-answer instances.
Stage 1: Generation of Candidate Options. Starting from CommonsenseQA questions and their gold answers, we sample 5,000 instances and use GPT-4o-mini to over-generate a diverse set of atomic answer candidates. The model is prompted to produce both plausible and implausible alternatives, with an emphasis on multi-step causal or situational reasoning rather than shallow lexical cues. This process yields 4-6 plausible and implausible candidate answers per question, spanning physical, social, and situational commonsense domains.
Stage 2: Refinement and Pruning. We then use GPT-4o-mini to refine and filter the generated candidates according to three criteria: (1) removing logically inconsistent or factually incorrect answers; (2) eliminating trivial options that can be resolved through keyword matching; and (3) identifying highly plausible options that satisfy most semantic and contextual constraints of the question but fail due to a non-obvious commonsense violation. This stage yields a curated set of three correct and four incorrect atomic options per question, each requiring multi-step inference.
Stage 3: Deterministic Logical Composition. In the final stage, we use a symbolic program to deterministically combine pairs of refined atomic options into logical compositions labeled with one of the three relations shown in Table 1. These relations are not strict propositional operators; instead, they encode commonsense judgments about joint plausibility. This procedure yields 14,997 instances in total, with 4,999 instances per base relation type.
The prompts for all stages of the construction pipeline are provided in Appendix A.3.
Human Validation
To ensure the commonsense validity of the refined atomic options, we conducted a human validation study using the awareness–consensus framework proposed by Whiting and Watts (2024). Two independent annotators evaluated a random sample of 250 questions (5% of Stage 2). For each atomic option, annotators assessed (1) whether they personally believed the option was correct (awareness) and (2) whether they believed most others would agree (consensus). This dual assessment captures both individual plausibility and perceived social agreement, aligning with recent views of commonsense as socially grounded. Disagreements between annotators were resolved through adjudication.
We report label accuracy on the adjudicated subset covering 50% of the question-level test split defined after Stage 2. Performance is high for AND (89.2%), OR (96.4%), and MIXED (88.4%), and reasonable for NEITHER/NOR (73.6%). These numbers reflect a reasonably accurate labeling of composed instances.
Inter-annotator agreement, measured using Cohen’s $\kappa$, indicated moderate agreement on human judgments ($\kappa=0.49$) before adjudication, suggesting that the refined options are both semantically coherent and socially interpretable.
Resulting Dataset
After Stage 2, we obtain 4,999 curated questions with refined atomic answer sets. In Stage 3, each question is expanded into logically structured instances evenly distributed across the three base relation types, with an additional Mixed setting in which operators are randomly assigned across options. After filtering, the final dataset contains 19,996 instances, with AND, OR, NEITHER/NOR, and MIXED each contributing 4,999 instances.
4 Task Formulation
Instances in Logical-CommonsenseQA consist of a question in natural language and four options. Unlike standard benchmarks where each option is an atomic answer, each option in our dataset is a logical composition of two independent atomic statements combined using one of three plausibility-level operators: AND, OR, or NEITHER/NOR (see Table  1).
Given a question, the model must select the option whose composite plausibility best matches the implied commonsense constraints. While the task remains a simple MCQ problem, it requires models to assess the plausibility of individual statements and reason about their interaction under the specified operator. We include both operator-specific conditions, in which all answer options for a question share the same logical operator, and a Mixed condition, in which different operators appear across options. While the operator-specific conditions allow for controlled evaluation of each relation type, the Mixed condition prevents models from exploiting operator-specific patterns and requires direct inference over the composed statements. Overall, this formulation preserves the MCQ format while substantially increasing the demands on compositional commonsense reasoning.
5 Experiments
We evaluate a diverse set of encoder-only classifiers, encoder–decoder models, and decoder-only LLMs on Logical-CommonsenseQA under zero-shot, few-shot, and supervised fine-tuning regimes. Full details on model selection, prompting strategies, and training and inference procedures are provided in Appendices A.1, A.2, and A.5.
Following the dataset construction described in Section 3, we use all 19,996 instances of Logical-CommonsenseQA  split into 11,996 training, 6,000 development, and 2,000 test examples. Splits are stratified by logical relation, ensuring an even 25% distribution across AND, OR, NEITHER/NOR, and Mixed conditions. We report accuracy and macro F1 as evaluation metrics.
For analysis, we further divide the test set into a human-validated (HV) subset and a non-validated (NV) subset, each comprising 50% of the test data. Table 2 summarizes the main results. Across all model families, two consistent patterns emerge: (1) all models perform reasonably well on AND and moderately on OR, and (2) performance collapses on NEITHER/NOR in zero-and few-shot settings, even for the strongest LLMs. Observed weaknesses are amplified in the Mixed condition, where different operators appear across options and models must implicitly infer the governing relation. F1 drops to 43–56% for all few-shot models, suggesting that these models do not reliably maintain operator-level representations and instead revert to surface-level heuristics. In contrast, fine-tuned models achieve 83–93% F1 across all operators, suggesting that the task is learnable with supervision and that zero- and few-shot failures, particularly on negation, reflect inference-time limitations rather than dataset artifacts.
Additional results—including varying numbers of shots, prompting strategies, and performance on the NV split—are reported in Appendix A.6. A comprehensive error analysis is also provided in Appendix A.7.
Comparison with CommonsenseQA. Finally, Table 3 underscores the central motivation for our benchmark. While LLaMA-3.1-8B achieves 72.2% accuracy on CommonsenseQA, its performance drops sharply on Logical-CommonsenseQA: 72% on AND, 62.2% on OR, 42.7% on MIXED, and only 13.9% on NEITHER/NOR. This discrepancy shows that benchmarks like CommonsenseQA substantially overestimate models’ commonsense reasoning ability by failing to test relational and compositional plausibility judgments.
Paradigm Model AND OR NN MIX
0-Shot LLaMA-3.3-70B 80.9 70.9 13.4 53.0
LLaMA-3.1-8B 71.9 62.2 13.1 41.8
Qwen2.5-7B 79.6 68.9 12.9 53.2
3-Shot LLaMA-3.3-70B 85.7 77.5 10.7 50.4
LLaMA-3.1-8B 71.3 61.3 6.1 42.8
Qwen2.5-7B 77.9 71.9 6.8 51.4
Fine-tuned Flan-T5-base 92.8 92.4 89.2 89.6
DeBERTa-v3-base 87.6 87.2 84.8 82.4
Table 2: Macro-F1 on the human-validated test set across plausibility relations. NN = NEITHER/NOR, MIX = MIXED
Dataset Acc F1
CommonsenseQA
LLaMA-3.1-8B 72.2 72.2
Logical-CommonsenseQA
AND 72.0 71.9
OR 62.2 62.2
NEITHER/NOR 13.9 13.1
MIXED 42.7 41.8
Table 3: 0-Shot performance of LLaMA-3.1-8B on the original CommonsenseQA versus each logical operator in Logical-CommonsenseQA.
6 Conclusions and Future Work
We introduce Logical-CommonsenseQA, a benchmark that reframes commonsense reasoning as selecting among logically composed answer options. By combining human-validated atomic statements with symbolic operators that encode conjunction, disjunction, and negation of plausibility, our work exposes reasoning failures obscured by traditional single-answer benchmarks. Experiments reveal that while LLMs handle conjunctive and, to a lesser extent, disjunctive reasoning, they fail sharply on negation-based composition, highlighting persistent gaps in plausibility evaluation and operator grounding. Future work includes extending our resource to generative settings and structured reasoning frameworks, expanding the operator set, scaling awareness-consensus validation, and studying transfer to open-ended QA, dialog, and planning tasks.
Limitations
Logical-CommonsenseQA offers a controlled framework for probing compositional commonsense reasoning, but several limitations remain. The logical operators (AND, OR and NEITHER/NOR) capture plausibility-level relations rather than full propositional semantics, and do not cover richer structures such as implication, exclusivity or temporal and causal reasoning. Extending the operator space would enable a more complete evaluation of symbolic and relational inference.
Model coverage is another limitation. Decoder-only LLMs are evaluated only via prompting, while smaller encoder-based models receive supervised training. A broader comparison across instruction-tuned and specialized architectures may clarify which components most improve logical composition.
Ethical Considerations
Logical-CommonsenseQA is derived from the publicly available CommonsenseQA dataset and consists of generic, hypothetical scenarios that do not reference identifiable individuals or personal contexts. To the best of our knowledge, the dataset does not contain private or sensitive information.
We used an LLM to initially generate atomic answer options and increase coverage and diversity. These candidates are subsequently refined and partially validated by human annotators using an awareness–consensus framework. As a result, the dataset reflects socially grounded plausibility judgments rather than objective ground truth, and may thus inherit cultural or contextual biases.
Appendix A Appendix
A.1 Model Selection
This section describes the architectures and evaluation settings of all models considered in our experiments, covering encoder-only, encoder-decoder, and decoder-only LLMs.
Encoder Models
We fine-tune DeBERTa-v3-base He et al. (2021), a pretrained encoder which encodes the concatenation of the question and one composite option and applies a classification head to score the four candidate answers.
Encoder–Decoder and Multi-Step Reasoning Models
We also fine-tune sequence-to-sequence models to generate the output label. Flan-T5-base Chung et al. (2024) is fine-tuned to output a single token corresponding to the chosen option. Additionally, we evaluate Entailer-11B Tafjord et al. (2022), a model designed to produce faithful chains-of-reasoning by recursively generating premises through backward-chaining and verifying them via self-querying. Although originally developed for faithful explanation generation, we use Entailer as a strong encoder-decoder baseline to examine whether reasoning-oriented models better capture the compositional structures in our dataset.
Decoder-Only LLMs
We evaluate several open-weight instruction-tuned LLMs in zero- and few-shot settings: Llama-3.3-70B-Instruct, Llama-3.1-8B-Instruct Dubey et al. (2024) and Qwen2.5-7B-Instruct Qwen et al. (2025). These models are not fine-tuned; instead, they are prompted to output a single capital letter corresponding to the predicted option. All models are evaluated on each logical subset defined in Table 1 as well as on the Mixed setting.
A.2 Prompting Setup
In this section, we describe the prompting strategy used for decoder-only models. For decoder-only LLMs, we adopt a unified multiple-choice prompting scheme. Each prompt presents the question, the four composite options and an instruction to answer with a single letter (A-D). We evaluate 0-, 1-, 2- and 3-shot settings using a fixed set of labeled examples.
Decoding uses temperature =0.0=0.0, top-p =0.9=0.9, and max_new_tokens=3=3. Outputs not corresponding to a valid answer option are treated as incorrect.
Chain-of-Thought (CoT)
We evaluate CoT prompting for Llama-3.1-8B-Instruct only, due to computational cost. CoT runs use a larger generation budget (max_new_tokens =200=200-300300).
A.3 Prompt Templates
This section provides the exact prompt templates used in our experiments. Prompts differ only in the number of in-context examples and the logical operator expressed in the answer options.
Zero-shot Prompt
In the zero-shot setting, we don’t provide any labeled examples.
Answer the following commonsense question by selecting the correct option. Respond with only the single capital letter (A, B, C, or D).
Question: <QUESTION>
A. <OPTION A>
B. <OPTION B>
C. <OPTION C>
D. <OPTION D>
Answer:
Few-shot Prompting
For few-shot (1-, 2-, and 3-shot) evaluation, we prepend labeled examples before the target question. The 1- and 2-shot prompts are strict prefixes of the 3-shot prompt, using the first one or two examples respectively.
For operator-specific subsets (AND, OR, NEITHER/NOR), all in-context examples match the operator being evaluated. For the Mixed setting, examples are sampled across operators.
Conjunctive Reasoning (AND) Prompt
Question: Sammy wanted to go to where the people were. Where might he go?
A. social venues AND quiet retreats
B. local events AND social venues
C. sports arenas AND quiet retreats
D. sports arenas AND train platforms
Answer: B
Question: The fox walked from the city into the forest, what was it looking for?
A. suitable prey AND shelter from disturbances
B. urban garden AND farm fields
C. natural habitat AND farm fields
D. suitable prey AND neighborhood pets
Answer: A
Question: What home entertainment equipment requires cable?
A. wireless speaker AND portable projector
B. television AND home theater system
C. wireless speaker AND gaming console
D. digital frame AND portable projector
Answer: B
Now answer the following commonsense question by selecting the correct option. Respond with only the single capital letter (A, B, C, or D).
Question: <TARGET AND QUESTION>
A. <OPTION A>
B. <OPTION B>
C. <OPTION C>
D. <OPTION D>
Answer:
Disjunctive Reasoning (OR) Prompt
Question: Sammy wanted to go to where the people were. Where might he go?
A. quiet retreats OR empty parks
B. local events OR empty parks
C. sports arenas OR empty parks
D. train platforms OR empty parks
Answer: B
Question: The forgotten leftovers had gotten quite old; he found it covered in mold in the back of his what?
A. living room OR dining area
B. utility drawer OR kitchen counter
C. living room OR utility drawer
D. cold storage OR utility drawer
Answer: D
Question: What do people use to absorb extra ink from a fountain pen?
A. absorbent paper OR notebook cover
B. ink storage OR writing surface
C. notebook cover OR writing surface
D. ink storage OR fabric swatch
Answer: A
Now answer the following commonsense question by selecting the correct option. Respond with only the single capital letter (A, B, C, or D).
Question: <TARGET OR QUESTION>
A. <OPTION A>
B. <OPTION B>
C. <OPTION C>
D. <OPTION D>
Answer:
Negation-Based Reasoning (NEITHER/NOR) Prompt
Question: Sammy wanted to go to where the people were. Where might he go?
A. NEITHER local events NOR train platforms
B. NEITHER sports arenas NOR train platforms
C. NEITHER social venues NOR sports arenas
D. NEITHER social venues NOR quiet retreats
Answer: B
Question: The only baggage the woman checked was a drawstring bag. Where was she heading with it?
A. NEITHER family gathering NOR local gathering
B. NEITHER airport travel NOR local gathering
C. NEITHER local gathering NOR short trip
D. NEITHER business engagement NOR local gathering
Answer: C
Question: The forgotten leftovers had gotten quite old, he found it covered in mold in the back of his what?
A. NEITHER cold storage NOR dining area
B. NEITHER sealed container NOR living room
C. NEITHER food cabinet NOR utility drawer
D. NEITHER living room NOR utility drawer
Answer: D
Now answer the following commonsense question by selecting the correct option. Respond with only the single capital letter (A, B, C, or D).
Question: <TARGET NEITHER/NOR QUESTION>
A. <OPTION A>
B. <OPTION B>
C. <OPTION C>
D. <OPTION D>
Answer:
Mixed Prompt
Question: The fox walked from the city into the forest, what was it looking for?
A. natural habitat AND shelter from disturbances
B. farm fields OR neighborhood pets
C. city park AND neighborhood pets
D. farm fields AND neighborhood pets
Answer: A
Question: Google Maps and other highway and street GPS services have replaced what?
A. historical navigation logs AND tourist information centers
B. NEITHER printed road maps NOR route planning software
C. manual navigation techniques AND tourist information centers
D. manual navigation techniques AND route planning software
Answer: D
Question: Where is a business restaurant likely to be located?
A. corporate office clusters AND at university campuses
B. business park entrances AND near conference venues
C. NEITHER near conference venues NOR at university campuses
D. NEITHER near conference venues NOR near tourist sites
Answer: B
Now answer the following commonsense question by selecting the correct option. Respond with only the single capital letter (A, B, C, or D).
Question: <TARGET MIXED QUESTION>
A. <OPTION A>
B. <OPTION B>
C. <OPTION C>
D. <OPTION D>
Answer:
Chain-of-Thought Prompting
For CoT evaluation, we add a brief instruction asking the model to reason about the plausibility of each atomic statement and how the logical operator applies before providing the final answer.
Carefully reason about each atomic statement and the logical operator. Then provide the final answer as a single capital letter (A–D).
Decoding Constraints
All decoder-only models are evaluated with near-deterministic decoding (temperature=0.0, top-p=0.9). Outputs are constrained to a single capital letter (A–D); any other output is treated as incorrect.
A.4 Annotation Guidelines
This section describes the guidelines provided to human annotators for validating refined atomic answer options in Stage 2 of dataset construction.
Annotation Task
Each annotation item consists of a commonsense question and a set of automatically labeled answer options. Options are pre-labeled as correct or incorrect by the refinement pipeline. Annotators are instructed to evaluate each option independently, without reconsidering or changing the question itself.
The goal of annotation is only verification: annotators judge whether the assigned labels are reasonable under commonsense reasoning.
Evaluation Criteria
For each option, annotators provide judgments along two dimensions, following the awareness–consensus framework of Whiting and Watts (2024):
  1. 1.
    Correctness (Awareness). Annotators indicate whether they personally believe the option is a valid answer to the question.
    • For options labeled as correct: “Do you believe this option is actually a valid answer?”
    • For options labeled as incorrect: “Do you believe this option is actually wrong or implausible?”
  2. 2.
    Commonsense (Consensus). Annotators indicate whether they believe most adults would agree with the judgment using everyday knowledge.
Each criterion is rated using a three-point scale: Yes, No, or Maybe. Annotators are required to provide a brief comment for any No or Maybe response (e.g., “could be true in some contexts” or “requires specialized knowledge”).
Definition of Commonsense
Annotators are instructed to interpret commonsense as knowledge that an average adult could reasonably be expected to possess, without relying on formal education, technical expertise, or domain-specific facts. Options requiring specialized scientific, historical, or technical knowledge are marked accordingly.
Handling Ambiguity
Annotators are encouraged to consider realistic context and pragmatic reasoning. When an option is plausible only under rare or contrived interpretations, it should be marked as Maybe rather than Yes. The goal is to assess typical human judgments, not theoretical possibility.
Adjudication
Two annotators independently evaluate each item. Disagreements, as well as any option receiving a Maybe are resolved through adjudication by a third reviewer. Only disputed cases are adjudicated; non-disputed options retain their original labels.
Illustrative Examples
Annotators are provided with worked examples illustrating borderline cases, such as partially plausible distractors, context-dependent interpretations, and options that are topically related but logically invalid. These examples are used to clarify expectations but are not treated as templates.
A.5 Training and Inference Details
This section reports the training configurations and inference settings used across all evaluated models.
Encoder-only and encoder-decoder models are fine-tuned with AdamW ($5\times 10^{-5}$ learning rate, weight decay $0.01$) for up to 10 epochs with early stopping (patience 3), using a maximum sequence length of 512 and an effective batch size of 8.
Decoder-only LLMs are evaluated without gradient updates under 0–3 shot prompting (with/without CoT) and outputs are constrained to a single label (A–D). Most experiments run on NVIDIA P100 GPUs. Only the largest models and CoT evaluations are run on A100 GPUs.
A.6 Additional Results
This section reports supplementary experimental results that complement the main findings discussed in Section 5. Table 4 reports accuracy and macro-F1 across all logical relations for human-validated (HV) and non-human-validated (NV) test subsets, including instruction-tuned and fine-tuned models. Table 5 presents macro-F1 for decoder-only LLMs under 0–3 shot prompting and CoT settings. Encoder-only and encoder–decoder models without fine-tuning exhibit uniformly low performance. Overall trends mirror the main paper: reasonable performance on AND and OR, and collapse on NEITHER/NOR across settings.
Model AND OR NN MIX
NV A NV F1 HV A HV F1 NV A NV F1 HV A HV F1 NV A NV F1 HV A HV F1 NV A NV F1 HV A HV F1
LLaMA-70B 76.4 76.2 80.8 80.9 72.4 72.3 71.2 70.9 13.6 13.4 14.0 13.4 53.7 53.4 53.2 53.0
LLaMA-8B 67.2 66.7 72.0 71.9 63.2 62.6 62.2 62.2 8.6 7.9 13.9 13.1 44.5 43.9 42.7 41.8
Qwen2.5-7B 76.2 76.2 79.6 79.6 70.4 70.5 68.8 68.9 13.6 11.8 13.3 12.9 54.0 53.7 53.2 53.2
Entailer-11B 5.6 5.5 3.2 3.2 20.8 20.8 18.4 18.4 22.0 22.0 24.0 24.1 16.8 16.8 22.4 22.3
Flan-T5-base 38.8 38.0 37.2 36.2 63.2 62.2 65.2 64.3 12.0 11.4 16.8 16.3 32.8 32.5 33.6 33.0
Flan-T5-base (FT) 94.8 94.8 92.8 92.8 97.6 97.6 92.4 92.4 93.2 93.2 89.2 89.2 87.2 87.2 89.6 89.6
DeBERTa-v3-base 28.8 28.6 27.2 27.2 23.6 23.8 23.2 23.1 30.8 30.9 27.2 27.4 28.8 28.8 28.4 28.4
DeBERTa-v3-base (FT) 90.8 90.8 87.6 87.6 88.0 88.0 87.2 87.2 89.2 89.2 84.8 84.8 84.8 84.8 82.4 82.4
Table 4: 0-shot vs. fine-tuned: Accuracy (A) and macro-F1 across four logical relations. NV = non-human-validated labels; HV = human-validated subset. NN = NEITHER/NOR, MIX = MIXED, FT = fine-tuned.
Model AND OR NN MIX
NV F1 HV F1 NV F1 HV F1 NV F1 HV F1 NV F1 HV F1
0-shot
LLaMA-70B 76.2 80.9 72.3 70.9 13.4 13.4 53.4 53.0
LLaMA-8B 66.7 71.9 62.6 62.2 7.9 13.1 43.9 41.8
Qwen-7B 76.2 79.6 70.5 68.9 11.8 12.9 53.7 53.2
1-shot
LLaMA-70B 80.3 84.8 76.9 79.5 7.9 11.3 54.7 53.1
LLaMA-8B 72.1 67.0 61.7 62.0 5.9 5.3 45.0 41.7
Qwen-7B 73.9 79.6 75.0 71.6 5.7 7.4 57.4 55.4
2-shot
LLaMA-70B 78.7 83.6 74.0 76.8 9.1 12.2 53.1 51.7
LLaMA-8B 71.4 75.4 61.6 63.9 6.0 6.6 46.1 41.2
Qwen-7B 72.3 77.5 70.6 69.6 5.7 6.9 55.2 52.0
3-shot
LLaMA-70B 80.0 85.7 75.1 77.5 7.8 10.7 52.4 50.4
LLaMA-8B 75.8 71.3 61.0 61.3 5.5 6.1 47.0 42.8
Qwen-7B 73.4 77.9 73.2 71.9 4.6 6.8 55.5 51.4
CoT (0-shot)
LLaMA-8B 65.3 67.9 63.0 62.5 10.7 8.2 45.8 42.7
Table 5: Macro-F1 across logical relations under 0–3 shot prompting and chain-of-thought (CoT). NV = non-human-validated labels; HV = human-validated subset. NN = NEITHER/NOR, MIX = MIXED.
A.7 Error Analysis
We analyze errors made by LLaMA-3.1-8B in the zero-shot setting. Table 6 summarizes representative failure patterns revealed by Logical-CommonsenseQA. Because each option is a compositional pair of atomic statements, errors expose not only mis-classifications but also deeper limitations highlighting how models fail to compose plausibility across logical operators.
Op Failure Pattern Representative Error (Gold-Pred)
AND Single-statement dominance Q: Heat applied to combustible materials?
Gold: emits toxic fumes AND decomposes into gases
Pred: ignites rapidly AND melts into liquid
OR Plausibility collapse Q: Heat applied to combustible materials?
Gold: ignites rapidly OR melts into liquid
Pred: melts into liquid OR burns without oxygen
NEITHER Negation inversion Q: Where do you buy a glass of wine?
Gold: NEITHER public park NOR home kitchen
Pred: NEITHER licensed restaurant NOR home kitchen
Plausibility dominance Q: Where might a child with no home go?
Gold: NEITHER school event NOR family gathering
Pred: NEITHER youth shelter NOR foster care
Table 6: Representative error patterns for LLaMA-3.1-8B (0-shot) across AND, OR, and NEITHER/NOR.
Conjunctive and Disjunctive Reasoning
Across both AND and OR settings, the dominant failure mode is the model’s tendency to recognize only a single plausible statement and ignore how plausibility must be composed across clauses. In AND questions, this appears as single-statement dominance: the model anchors on a highly plausible clause and fails to reject an implausible partner. For example, in the combustibility question, it predicts ignites rapidly AND melts into liquid, correctly identifying ignition while overlooking that combustible materials do not melt.
The same underlying behavior manifests in OR questions. Rather than exploiting the requirement that at least one clause be plausible, the model treats OR as indicating thematic similarity. In the combustibility example, it predicts melts into liquid OR burns without oxygen, selecting two implausible but fire-related statements. These errors indicate that models do not enforce the compositional constraint imposed by the operator; instead, they rely on surface-level plausibility cues and semantic similarity.
Negation-Based Reasoning
NEITHER/NOR proves most challenging. We observe negation inversion, where the model selects the most plausible pair of statements despite the operator requiring both to be implausible. For example, in the wine-purchase question, the model predicts NEITHER licensed restaurant NOR home kitchen, incorrectly negating the true location. A second pattern, plausibility dominance, reflects the model’s tendency to retain highly plausible statements even under negation. For example, in the homeless-child example, the model selects NEITHER youth shelter NOR foster care, two locations that are plausible options. These behaviors indicate a broader difficulty with composing negation and plausibility simultaneously.

### v2

@@
 This procedure yields 14,997 instances in total, with 4,999 instances per base relation type.
-The prompts for all stages of the construction pipeline are provided in Appendix A.3.
+Prompt templates for all stages of the construction pipeline are provided in Appendix A.3.1.
 Human Validation
@@
 A.3 Prompt Templates
-This section provides the exact prompt templates used in our experiments. Prompts differ only in the number of in-context examples and the logical operator expressed in the answer options.
+This section provides the prompt templates used for dataset construction and evaluation.
+A.3.1 Dataset Construction Prompts
+We describe the prompts used in each stage of dataset construction below.
+Stage 1: Candidate Option Generation
+Given a commonsense question and its original correct and incorrect answers, generate additional atomic answer options that require non-trivial reasoning.
+Instructions:
+  • Generate 4–6 additional correct options that are independently plausible and require multi-step commonsense reasoning (e.g., causal, situational, or social inference).
+  • Generate 4–6 incorrect options that are highly plausible and contextually related to the question, but fail due to a subtle violation (e.g., incorrect purpose, mismatched context, or missing constraint).
+  • Avoid trivial answers, obvious domain mismatches, or options that can be rejected through surface-level keyword matching.
+  • Ensure each option can be evaluated independently, without relying on other options or comparisons.
+Output: A JSON object containing expanded lists of correct and incorrect atomic answer options.
+Stage 2: Option Refinement and Pruning
+Given a commonsense question and a set of generated atomic answer options, refine and filter the options to produce a challenging and internally consistent reasoning set.
+Instructions:
+  • Remove options that are logically inconsistent, factually incorrect, trivial, or mismatched to the question category.
+  • Refine correct options to eliminate direct or overly obvious answers while preserving plausibility and the need for multi-step inference.
+  • Refine incorrect options into near-miss distractors that satisfy most semantic and contextual constraints but fail on exactly one subtle causal, temporal, or pragmatic nuance.
+  • Enforce consistency in option length and prohibit explanatory or justificatory phrasing.
+Output: A JSON object containing a fixed set of refined correct and incorrect atomic options.
+Stage 3: Logical Composition
+In Stage 3, refined atomic options are combined using deterministic symbolic rules to form composed answer options labeled with logical relations.
+This stage does not involve prompting or any additional model inference.
+A.3.2 Evaluation Prompts
+Prompts differ only in the number of in-context examples and the logical operator expressed in the answer options.
 Zero-shot Prompt